# PRISM: Leveraging Prototype Patient Representations with Feature-Missing-Aware Calibration for EHR Data Sparsity Mitigation

### v5

@@ 5.1 Benchmarked Real-World Datasets
Table 2: Statistics of datasets after preprocessing. The proportion demonstrates the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
@@
5.2 Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets [29, 30] as shown in our selected datasets.
5.3 Baseline Models
@@
5.3.1 Imputation-based Methods
@@
  • MICE [13] addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model [31].
  • GRU-D [11] incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
5.3.2 EHR-specific Models
@@
  • RETAIN [1] is a hierarchical attention-based interpretable model. It attends the EHR data in a reverse time order so that recent clinical visits are likely to receive higher attention.
  • AdaCare [16] is a GRU-based network that utilizes a multi-scale dilated convolutional module to capture the long and short-term historical variation.
  • ConCare [2] utilizes multi-channel GRU with a time-aware attention mechanism to extract clinical features and re-encode the clinical information by capturing the interdependencies between features.
  • GRASP [3] is a generic framework for healthcare models, which leverages the information extracted from patients with similar conditions to enhance the cohort representation learning results.
  • M3Care [12] resolves the missing modality issue in EHR data by utilizing similar patients’ existing modalities. However, it does not address the issue of missing features within available modalities.
  • SAFARI [5] learns patient health representations by applying a clinical-fact-inspired, task-agnostic correlational sparsity prior to medical feature correlations, using a bi-level optimization process that involves both inter- and intra-group correlations.
  • AICare [17] also includes a multi-channel feature extraction module and an adaptive feature importance recalibration module. It learns personalized health status embeddings with static and dynamic features.
5.3.3 Ablation Models
@@
5.4 Implementation Details
5.4.1 Hardware and Software Configuration
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s system memory (RAM) size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 [32], PyTorch Lightning 2.0.5 [33], and pyehr [34].
5.4.2 Model Training and Hyperparameters
AdamW [35] is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in AICare [17].
6 Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets. The dataset preprocessing scripts, developed model’s code and all baseline models’ code are attached at the Supplementary Material for reproducibility.
6.1 Experimental Results
Table 3: Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@ 6.1 Experimental Results
6.2 Ablation Study
6.2.1 Comparing with Reduced Versions
@@
6.2.2 Comparing with Internal Components
@@
Table 4: Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@
6.3 Observations and Analysis
6.3.1 Robustness to Data Sparsity
@@
Figure 2: AUPRC performance across 5 sparsity levels in MIMIC-IV in-hospital mortality prediction task. PRISM significantly outperforms other models in extremely sparse scenarios on both sparsity settings.
6.3.2 Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
@@
Figure 3: AUPRC and F1 score performance on various prototype patient cohort size in MIMIC-IV in-hospital mortality prediction task. With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
6.3.3 Variations of Similarity Measures
@@
Table 5: Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@
6.3.4 Model Efficiency and Complexity
We evaluate the efficiency and complexity of PRISM against SOTA EHR-specific models baseline models. Table 6 compares the total parameter count and runtime of PRISM with these models on the MIMIC-IV dataset, using the same hidden dimension (128) and batch size (1024) hyperparameters. PRISM achieves a competitive balance between a low parameter count (215K) and an efficient runtime (69.03s).
Table 6: Computational comparison with latest SOTA EHR-specific models on the MIMIC-IV dataset. All models employ the same 128 hidden dimension and 1024 batch size hyperparameters.
Model # Parameter Runtime (5 epochs)
ConCare (AAAI-2020) 3.3M 96.62s
GRASP (AAAI-2021) 3.3M 97.68s
M3Care (KDD-2022) 469K 68.29s
SAFARI (TKDE-2022) 1.8M 78.97s
AICare (Patterns-2023) 2.2M 72.97s
PRISM 215K 69.03s
Table 7 presents the data preparation time for PRISM on MIMIC-III and MIMIC-IV datasets, including preprocessing, training, validation, and testing. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the entire pipeline can be completed within 10 minutes for each dataset.
Table 7: Data preparation time comparison. Prep. denotes Preprocessing, Val. denotes Validation. All with 50 epochs of training, validate at the end of each epoch. The units are seconds.
@@ 6.3.4 Model Efficiency and Complexity
6.3.5 Cross-Feature Attention Map
@@
Figure 4: Cross-feature attention maps from PRISM: Without (Left) / With (Right) feature confidence calibration. The maps use data from a single MIMIC-IV patient record to show PRISM’s reduction in attention to unreliable features (capillary refill rate, fraction inspired oxygen, height) due to high missing rates.
6.3.6 Patient Representation Visualization
@@
Figure 5: t-SNE visualization of patient representations from PRISM. (a) shows the embeddings without the missing-feature-calibration process, and (b) depicts embeddings with the process. (b)’s representations are more compact among dead outcome patients, showcasing it learns better representations.
6.3.7 Feature Decay Rates Observation
@@
Figure 6: Adaptive learning feature decay rates. The graph shows varying decay rates: high for acute-indicator features like heart rate and pH, and low for longer-term relevant features like sex and systolic blood pressure.
7 Limitations and Further Work
We propose potential limitations of our current model and suggest directions for future research:
  • Fairness Concerns: Like many machine learning applications, it is essential to evaluate our model’s fairness across various demographic groups [7]. Discovering bias issue in similar patients cohort could be further explored.
  • Scalability Issues: The proposed model’s scalability to more massive datasets or integration within real-time healthcare systems is yet to be assessed.
  • Prototype Patient Representations: Though leveraging prototype patient representations has shown promise, further work is needed to understand the diversity of these prototypes. More intricate mechanisms for prototype generation, beyond similarity metrics, could also be explored.
8 Conclusions
@@
This study, which involves analyzing Electronic Health Records (EHR) from the MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets, is dedicated to maintaining high ethical standards. All datasets used are de-identified to ensure patient confidentiality and privacy. By signing the data use agreement, we reaffirm our commitment to responsible data handling and usage. Our methodology is designed to minimize potential harm and to produce findings that are as unbiased and equitable as possible, reflecting the diverse and complex nature of medical data. We have also taken rigorous steps to ensure that our research adheres to these ethical values.

### v6

@@ 5.1. Benchmarked Real-World Datasets
Table 2. Statistics of datasets after preprocessing. The proportion demonstrates the percentage of the label with value $1$. $Out.$ denotes Mortality Outcome, $Re.$ denotes Readmission.
@@
5.2. Evaluation Metrics
We assess the binary classification performance using AUROC, AUPRC and F1. Here we emphasize AUPRC as the main metric due to it is informative when dealing with highly imbalanced and skewed datasets (Kim and Hwang, 2022; Davis and Goadrich, 2006) as shown in our selected datasets.
5.3. Baseline Models
@@
5.3.1. Imputation-based Methods
@@
  • MICE (Van Buuren and Groothuis-Oudshoorn, 2011) addresses missing data in EHR through iterative imputation, with subsequent analysis using an LSTM model (Hochreiter and Schmidhuber, 1997).
  • GRU-D (Che et al., 2018) incoporates both the last observed and global mean values in the GRU network. Additionally, GRU-D utilizes an exponential decay mechanism to manage the temporal dynamics of missing values.
5.3.2. EHR-specific Models
@@
  • RETAIN (Choi et al., 2016) is a hierarchical attention-based interpretable model. It attends the EHR data in a reverse time order so that recent clinical visits are likely to receive higher attention.
  • AdaCare (Ma et al., 2020a) is a GRU-based network that utilizes a multi-scale dilated convolutional module to capture the long and short-term historical variation.
  • ConCare (Ma et al., 2020b) utilizes multi-channel GRU with a time-aware attention mechanism to extract clinical features and re-encode the clinical information by capturing the interdependencies between features.
  • GRASP (Zhang et al., 2021) is a generic framework for healthcare models, which leverages the information extracted from patients with similar conditions to enhance the cohort representation learning results.
  • M3Care (Zhang et al., 2022) resolves the missing modality issue in EHR data by utilizing similar patients’ existing modalities. However, it does not address the issue of missing features within available modalities.
  • SAFARI (Ma et al., 2022) learns patient health representations by applying a clinical-fact-inspired, task-agnostic correlational sparsity prior to medical feature correlations, using a bi-level optimization process that involves both inter- and intra-group correlations.
  • AICare (Ma et al., 2023) also includes a multi-channel feature extraction module and an adaptive feature importance recalibration module. It learns personalized health status embeddings with static and dynamic features.
5.3.3. Ablation Models
@@
5.4. Implementation Details
5.4.1. Hardware and Software Configuration
All runs are trained on a single Nvidia RTX 3090 GPU with CUDA 11.8. The server’s system memory (RAM) size is 64GB. We implement the model in Python 3.11.4, PyTorch 2.0.1 (Paszke et al., 2019), PyTorch Lightning 2.0.5 (Falcon, 2019), and pyehr (Zhu et al., 2024).
5.4.2. Model Training and Hyperparameters
AdamW (Loshchilov and Hutter, 2017) is employed with a batch size of 1024 patients. All models are trained for 50 epochs with an early stopping strategy based on AUPRC after 10 epochs without improvement. The learning rate ${0.01,0.001,0.0001}$ and hidden dimensions ${64,128}$ are tuned using a grid search strategy on the validation set. The searched hyperparameter for PRISM is: 128 hidden dimensions, 0.001 learning rate, and 256 prototype patient cluster numbers. Performance is reported in the form of mean±std of 5 runs with random seeds ${0,1,2,3,4}$ for MIMIC-III and MIMIC-IV datasets, and apply bootstrapping on all test set samples 10 times for the Challenge-2012 and eICU datasets, following practices in Ma et al. (2023).
Table 3. Benchmarking results on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets. Bold indicates the best performance. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@ 5.4.2. Model Training and Hyperparameters
6. Experimental Results and Analysis
We conduct the in-hospital mortality and 30-day readmission prediction task on MIMIC-III and MIMIC-IV datasets, in-hospital mortality prediction task on Challenge-2012 and eICU datasets.
6.1. Experimental Results
6.2. Ablation Study
6.2.1. Comparing with Reduced Versions
@@
6.2.2. Comparing with Internal Components
@@
Table 4. Performance comparison of internal components on the MIMIC-IV mortality prediction task. “Feat. Conf.” means “Feature Confidence” and “Sim. Meas.” denotes “Similarity Measure”. Bold denotes the best performance within each components, Red denotes the highest performance among all comparisons. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@
6.3. Observations and Analysis
6.3.1. Robustness to Data Sparsity
@@
Figure 2. AUPRC performance across 5 sparsity levels in MIMIC-IV in-hospital mortality prediction task. PRISM significantly outperforms other models in extremely sparse scenarios on both sparsity settings.
6.3.2. Sensitiveness to Cohort Size and Effectiveness of Missing-Feature-Aware Module
@@
Figure 3. AUPRC and F1 score performance on various prototype patient cohort size in MIMIC-IV in-hospital mortality prediction task. With the missing-feature-aware module, PRISM outperforms its counterpart without. It also shows PRISM is not sensitive to the cohort size.
6.3.3. Variations of Similarity Measures
@@
Table 5. Comparing different similarity measures in MIMIC-IV in-hospital mortality prediction task. Performance is reported in the form of mean±std. All metric scores are multiplied by 100 for readability purposes.
@@
6.3.4. Model Efficiency and Complexity
We evaluate the efficiency and complexity of PRISM in terms of parameter count, runtime, and data preparation time. PRISM achieves a competitive balance between a low parameter count (215K) and an efficient runtime (69.03s for 5 epochs) on the MIMIC-IV dataset, using a hidden dimension of 128 and batch size of 1024. Table 6 presents the data preparation time for PRISM on MIMIC-III, MIMIC-IV, Challenge-2012, and eICU datasets, including preprocessing, training, validation, and testing. The PRISM pipeline seamlessly computes feature missing statuses in the LOCF pipeline with little extra cost. Moreover, the entire pipeline can be completed within 10 minutes for each dataset.
Table 6. Data preparation time comparison. Prep. denotes Preprocessing, Val. denotes Validation. All with 50 epochs of training, validate at the end of each epoch. The units are seconds.
@@ 6.3.4. Model Efficiency and Complexity
6.3.5. Cross-Feature Attention Map
@@
Figure 4. Cross-feature attention maps from PRISM: Without (Left) / With (Right) feature confidence calibration. The maps use data from a single MIMIC-IV patient record to show PRISM’s reduction in attention to unreliable features (capillary refill rate, fraction inspired oxygen, height) due to high missing rates.
6.3.6. Patient Representation Visualization
@@
Figure 5. t-SNE visualization of patient representations from PRISM. (a) shows the embeddings without the missing-feature-calibration process, and (b) depicts embeddings with the process. (b)’s representations are more compact among dead outcome patients, showcasing it learns better representations.
6.3.7. Feature Decay Rates Observation
@@
Figure 6. Adaptive learning feature decay rates. The graph shows varying decay rates: high for acute-indicator features like heart rate and pH, and low for longer-term relevant features like sex and systolic blood pressure.
7. Limitations and Further Work
We identify key limitations and future research directions:
  • Fairness Concerns: Evaluate the model’s fairness across various demographic groups and explore bias in similar patient cohorts.
  • Scalability Issues: Assess the scalability of the proposed model to larger datasets or its integration within real-time healthcare systems.
  • Prototype Patient Representations: Understand the diversity of prototype patient representations and explore more intricate mechanisms for prototype generation beyond similarity metrics.
8. Conclusions
@@
This study analyzes de-identified EHR data from MIMIC-III, MIMIC-IV, PhysioNet Challenge 2012, and eICU datasets. We adhere to data use agreements, prioritize patient privacy, and strive for unbiased, equitable findings that reflect the complexity of medical data. Our methodology aims to minimize potential harm and uphold ethical research standards.
Acknowledgements.
This work was supported by the National Science and Technology Major Project(2022ZD0116401), the National Natural Science Foundation of China (U23A20468), and Xuzhou Scientific Technological Projects (KC23143).